A point is equidistant from points A(0,0) B(4,0)and C (4,4). Its coordinates are _____.
A (2,1)
B (1,3)
c (0,4)
d (2,2)

Answer by richwmiller(17219)

point (2,2) is 2sqrt2 from those three points
d=sqrt((x2-x1)^2 +(y2-y1)^2)
sqrt((4-2)^2 +(4-2)^2)=sqrt(4+4)=sqrt8=2sqrt2
sqrt((4-2)^2 +(0-2)^2)=sqrt(4+4)=sqrt8=2sqrt2
sqrt((0-2)^2 +(0-2)^2)=sqrt(4+4)=sqrt8=2sqrt2